In 2015, it was reported that 22% of smartphones in operation in the United States were iPhones. If this is still accurate, what is the probability that less than 2 of the next 7 smartphone users you see have iPhones?
Answer #1

Binomial distribution: P(x) = nCx px qn-x

P(uses iPhone), p = 0.22

q = 1 - 0.22 = 0.78

Sample size, n = 7

P(less than 2 of the next 7 smartphone users you see have iPhones) = P(x < 2)

= P(0) + P(1)

= 0.787 + 7x0.22x0.786

= 0.52246